5. Hydrogen gas has a nonpolluting combustion product (water vapor). It is used as a fuel aboard the space shuttle and in earthbound cars with prototype engines: 2H20 g) O2(g) a. Express the rate in terms of changes in [Hal, CO2, and [H20] with time. b. When CO2] is decreasing at O.23 moyL. s, at what rate is CH20] increasing?
Answer #1

a)

rate = -(1/2)d[H2]/dt

rate = - d[O2]/dt

rate = (1/2)d[H2O]/dt

b)

we have,

rate = - d[O2]/dt = (1/2)d[H2O]/dt

- d[O2]/dt = (1/2)d[H2O]/dt

- (-0.23 mol/L.s) = (1/2)*d[H2O]/dt

d[H2O]/dt = 0.46 mol/L.s

Answer:

H2O is increasing at the rate of 0.46 mol/L.s